OpenStudy (anonymous):

Find the slope of the line that goes through the given points. (-1, 4), (5, 4)

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (-1,4) x1 = -1 and y1 = 4 (5,4) x2 = 5 and y2 = 4 now sub slope = (4 - 4) / (5 - (-1) you can finish with the math... keep in mind, horizontal lines have 0 slope and vertical lines have undefined slopes.